# Dynamics of inviscid annular supersonic jets

### v1

@@ 5 Conclusions and perspectives
The linear dynamics of inviscid annular supersonic jets, similar to those that are encountered in the exhaust a converging diverging nozzle, was explored in this article. The aim was to provide insights on the physical origins of tonal dynamics, observed in very limited expansion regimes in experiments (Jaunet et al., 2017; Baars et al., 2012b), and possibly responsible of unsteady side-loads (Bakulu et al., 2021; Martelli et al., 2020). The focus was therefore put to the first azimuthal Fourier mode of flow fluctuations, the only ones responsible for off-axis loads in axisymmetric nozzles. The dynamical properties of such fluctuations were computed on base flows with varying inner mixing layer radial position or varying mixing layer thicknesses. Finally, the effect of a variable expansion regime, i.e. increase in NPR as the flow can encounter during the start-up of an engine, was explored.
@@
As expected from the shape of the base flows, two unstable Kelvin-Helmholtz were found in all cases. The KH wave supported by the outter mixing layer has most of the time been found more unstable than the inner one. This is coherent with the differences in velocity gradients in the two mixing layers. Each of the associated eigenfunctions shows a maximum at the mixing layers location, typical of the KH wave. As could be seen, the outter KH waves eigenfunctions never showed tremendous changes while the flow parameters were varied. In the contrary, the inner KH eigenfunctions has shown to be more affected by base flow changes. Their structure can show a secondary peak at the location of the outter mixing layer and, more importantly, they seem to have a more pronounced signature outside of the jet than the outter KH wave. We suppose that this can be attributed to the fact that the inner and outter mixing layer have velocity gradient of opposite sign, providing different decay rates for both waves. The fact that the inner KH wave has support outside of the jet must be noticed as it allows this wave to exchange energy with external acoustics or upstream travelling waves. It underpins the fact that the inner mixing layer, emanating from the Mach disk triple point, is a probable source of instability in the resonance observed in convergent-divergent nozzles, as conjectured in Jaunet et al. (2017) and observed numerically by Bakulu et al. (2021).
@@
The article purposely took interest in describing the guided-jet modes structure and dynamical characteristics as they have recently shown responsible for the feedback process in jet resonances (see Edgington-Mitchell et al. (2022); Varé & Bogey (2022) for recent references). The most striking result of the current study is the robustness of the guided-jet modes: they were found in all cases and their eigenfunctions were shown to very poorly sensitive to the base flow parameters tested. This is clear indication that whenever a jet flow is considered one must pay attention to these waves as they could be involved in the dynamics. In our case, the dispersion relation and the existence of the upstream-travelling wave depends strongly on the inner mixing layer position inside the jet but rather poorly with the mixing layer thicknesses The latter results suggesting that they could be observed rather far downstream in an exhausting jet that diffuses in the surrounding.
@@
Finally, a simplified base flow model was derived in order to evaluate the stability properties of the flow with varying NPR but arbitrarily fixed mixing layer thicknesses. Like the hypothetical base flows used in the beginning of the study, two Kelvin-Helmholtz modes and numerous guided jet modes could be identified in the eigenspectra of the analysis, with characteristics found in line with the previous results. Interestingly, we have observed that the inner KH wave has rather small growth rate at low Mach number (NPR) and suddenly increases above the Mach number at which resonances was experimentally observed. Although this can be fortuitous, this yet another indication that we have pinpointed the correct waves at play in the resonances in such flows and that the inner mixing layer is of great importance in the dynamics of such flows.
@@
Although the results of this study revealed interesting feature of the dynamics of annular supersonic jets, we must recall that an important number of simplifications were made in order to be able to conduct the analysis. These assumptions are very likely the reason why the prediction of resonance frequency, as was done in the screeching jet case, not possible. This provides obvious path for improvement in both the base flow modeling, by considering a more representative control volume (Li & Ben-Dor, 1998), or even the dynamical model via the consideration of the influence of the nozzle walls for example.

### v2

@@ 4 Conclusions and perspectives
The linear dynamics of inviscid annular supersonic jets, similar to those encountered in the exhaust of a converging diverging nozzle, was explored in this article. The aim was to provide insights on the physical origins of tonal dynamics, observed in very limited expansion regimes in experiments (Jaunet et al., 2017; Baars et al., 2012b), and possibly responsible of unsteady side-loads (Bakulu et al., 2021; Martelli et al., 2020). The focus was therefore put on the first azimuthal Fourier mode of flow fluctuations, the only one responsible for off-axis loads in axisymmetric nozzles. The dynamical properties of such fluctuations were computed on base flows with varying inner mixing layer radial position or varying mixing layer thicknesses. Finally, the effect of a variable expansion regime, i.e. increase in NPR as the flow can encounter during the start-up of an engine, was explored.
@@
As expected from the shape of the base flows, two unstable Kelvin-Helmholtz were found in all cases. The KH wave supported by the outer mixing layer has most of the time been found more unstable than the inner one. This is coherent with the differences in velocity gradients in the two mixing layers. Each of the associated eigenfunctions shows a maximum at the mixing layers location, typical of the KH wave. The outer KH waves eigenfunctions never showed significant changes as function of the variation of flow parameters. On the contrary, the inner KH eigenfunctions have shown to be more affected by base flow changes. Their structure can show a secondary peak at the location of the outer mixing layer and, more importantly, they seem to have a more pronounced signature outside of the jet than the outer KH wave. The fact that the inner KH wave has support outside of the jet must be noticed as it allows this wave to exchange energy with external acoustics or upstream travelling waves. It underpins the fact that the inner mixing layer, emanating from the Mach disk triple point, is a probable source of instability in the resonance observed in convergent-divergent nozzles, as conjectured in Jaunet et al. (2017) and observed numerically by Bakulu et al. (2021).
@@
The article purposely took interest in describing the guided-jet modes structure and dynamical characteristics as they have recently shown responsible for the feedback process in jet resonances (see Edgington-Mitchell et al. (2022); Varé & Bogey (2022) for recent references). The most striking result of the current study is the robustness of the guided-jet modes: they were found in all cases and their eigenfunctions were shown to be very poorly sensitive to the base flow parameters tested. This is a clear indication that whenever a jet flow is considered one must pay attention to these waves as they could be involved in the dynamics. In our case, the dispersion relation and the existence of the upstream-travelling wave depends strongly on the inner mixing layer position inside the jet but rather poorly with the mixing layer thicknesses The latter results suggesting that they could be observed rather far downstream in an exhausting jet that diffuses in the surrounding.
@@
Finally, a simplified base flow model was derived in order to evaluate the stability properties of the flow with varying NPR but arbitrarily fixed mixing layer thicknesses. Like the hypothetical base flows used in the beginning of the study, two Kelvin-Helmholtz modes and numerous guided jet modes could be identified in the eigenspectra of the analysis, with characteristics found in line with the previous results. Interestingly, we have observed that the inner KH wave has rather small growth rate at low Mach number (NPR) and suddenly increases above the Mach number at which resonances was experimentally observed. Although this can be fortuitous, this is yet another indication that we have possibly pinpointed the correct waves at play in the resonances in such flows and that the inner mixing layer is of great importance in the dynamics of such flows.
@@
Although the results of this study revealed interesting feature of the dynamics of annular supersonic jets, we must recall that an important number of simplifications were made in order to be able to conduct the analysis. These assumptions are very likely the reason why the prediction of resonance frequency, as was done in the screeching jet case, was not possible. This provides obvious path for improvement in both the base flow modeling, by considering a more representative control volume (Li & Ben-Dor, 1998), the consideration of the influence of the nozzle walls or even the relaxation of the parallel flow assumption by the use of a global stability analysis.
Appendix
4.1 Mass conservation
Applying mass conservation on a control volume delimited by the nozzle throat and the attached flow, as presented in grey in figure3, we can write:
$\rho_{*}U_{*}D_{c}^{2}$ $=$ $\rho_{i}U_{i}D_{i}^{2}+\rho_{a}U_{a}(D_{e}^{2}-D_{i}^{2}),$ (16)
where $(\cdot)_{*}$ denotes sonic variables at the throat.
Using perfect gaz relations and introducing the speed of sound, such that $U=aM=\sqrt{\gamma rT}\cdot M$ and recalling that $M_{*}=1$, we have :
$\frac{P^{*}}{rT^{*}}a^{*}M^{*}{D_{c}}^{2}$ $=$ $\frac{P_{i}}{rT_{i}}a_{i}M_{i}D_{i}^{2}+\frac{P_{i}}{rT_{i}}a_{i}M_{i}(D_{e}^{2}-D_{i}^{2})$
$\frac{P^{*}}{\sqrt{rT^{*}}}{D_{c}}^{2}$ $=$ $\frac{P_{i}}{\sqrt{rT_{i}}}M_{i}D_{i}^{2}+\frac{P_{a}}{\sqrt{rT_{a}}}M_{a}(D_{e}^{2}-D_{i}^{2})$
Then using isentropic relation to link static and total variables, we obtain :
${\left(1+\frac{\gamma-1}{2}\right)^{-\frac{\gamma+1}{2(\gamma-1)}}}\frac{P_{t0}}{\sqrt{T_{t0}}}D_{c}^{2}$ $=$ $\frac{P_{i}}{\sqrt{T_{i}}}\mathcal{M}_{i}D_{i}^{2}+\frac{P_{a}}{\sqrt{T_{a}}}\mathcal{M}_{a}(1-D_{i}^{2})$
Choosing $P_{a}=P_{i}=P_{\infty}$, we can further simplify the mass conservation equation:
${\left(1+\frac{\gamma-1}{2}\right)^{-\frac{\gamma+1}{2(\gamma-1)}}}\frac{P_{t0}}{P_{\infty}}{D_{c}^{2}}$ $=$ $\frac{\sqrt{T_{t0}}}{\sqrt{T_{i}}}\mathcal{M}_{i}\frac{D_{i}^{2}}{+}\frac{\sqrt{T_{t0}}}{\sqrt{T_{a}}}\mathcal{M}_{a}\frac{(D_{e}^{2}-D_{i}^{2})}{}$
${\left(1+\frac{\gamma-1}{2}\right)^{-\frac{\gamma+1}{2(\gamma-1)}}}\frac{P_{t0}}{P_{\infty}}{D_{c}^{2}}$ $=$ ${D_{i}^{2}}{\mathcal{M}_{i}\sqrt{1+\frac{\gamma-1}{2}\mathcal{M}_{i}^{2}}}$
$+{(D_{e}^{2}-D_{i}^{2})}{\mathcal{M}_{a}\sqrt{1+\frac{\gamma-1}{2}\mathcal{M}_{a}^{2}}},$
which can be written in the more compact form presented in the document :
$\beta_{m}\frac{D_{c}^{2}}{D_{e}^{2}}$ $=$ $(\mu_{i}-\mu_{a})\frac{D_{i}^{2}}{D_{e}^{2}}+\mu_{a}$ (17)
where:
$\mu_{i,a}$ $=$ ${\mathcal{M}_{i,a}\sqrt{1+\frac{\gamma-1}{2}\mathcal{M}_{i,a}^{2}}}$
4.2 Momentum conservation
Now considering the momentum conservation along the axial direction on the same control volume, and neglecting body and viscous forces, leads to:
$\gamma P_{*}{\mathcal{M}_{*}}^{2}D_{c}^{2}+P_{*}D_{c}^{2}$ $=$ $\gamma P_{i}\mathcal{M}_{i}^{2}D_{i}^{2}+P_{i}D_{i}^{2}+\gamma P_{a}\mathcal{M}_{a}^{2}(D_{e}^{2}-D_{i}^{2})$
$+P_{a}(D_{e}^{2}-D_{i}^{2})-F_{w}$
where $F_{w}$ represents the pressure forces acting on the nozzle wall. Using isentropic relations, isobaric assumption (as for the mass conservation) and rearanging provides:
$(1+\gamma)\left(1+\frac{\gamma-1}{2}\right)^{\frac{-\gamma}{\gamma-1}}\frac{P_{t0}}{P_{\infty}}D_{c}^{2}-\frac{F_{w}}{P_{\infty}}$ $=$ $D_{i}^{2}(\gamma\mathcal{M}_{i}^{2}+1)+D_{e}^{2}(\gamma\mathcal{M}_{a}^{2}+1)-D_{i}^{2}(\gamma\mathcal{M}_{a}^{2}+1),$
which can be written in the more compact form:
${\beta_{q}}\frac{D_{c}^{2}}{D_{e}^{2}}-\frac{F_{w}}{P_{\infty}}$ $=$ $\gamma\frac{D_{i}^{2}}{D_{e}^{2}}\left(\mathcal{M}_{i}^{2}-\mathcal{M}_{a}^{2}\right)+\left(\gamma\mathcal{M}_{a}^{2}+1\right),$ (18)
where:
$\beta_{q}$ $=$ $(1+\gamma)\left(1+\frac{\gamma-1}{2}\right)^{\frac{-\gamma}{\gamma-1}}\frac{P_{t0}}{P_{\infty}},$
From the latter equation we can get an analytical expression for $D_{i}^{2}$:
$D_{i}^{2}$ $=$ $D_{e}^{2}\frac{1}{\gamma\left(\mathcal{M}_{i}^{2}-\mathcal{M}_{a}^{2}\right)}\left[{\beta_{q}}D_{c}^{2}-\frac{F_{w}}{P_{\infty}}-\left(\gamma\mathcal{M}_{a}^{2}+1\right)\right],$
which can be plugged into the mass conservation equation leading to the equation for the annular Mach number presented in the document:
${\frac{\mu_{i}-\mu_{a}}{\gamma\left(\mathcal{M}_{i}^{2}-\mathcal{M}_{a}^{2}\right)}\left[{\beta_{q}}{D_{c}^{2}}-\frac{F_{w}}{P_{\infty}}-\left(\gamma\mathcal{M}_{a}^{2}+1\right)\right]}+\mu_{a}-\beta_{m}$ $=$ 00 0 (19)
where:
$\mu_{i,a}$ $=$ ${\mathcal{M}_{i,a}\sqrt{1+\frac{\gamma-1}{2}\mathcal{M}_{i,a}^{2}}}$